OpenStudy (cutestgirlaround):

Two cars traveled equal distances in different amounts of time. Car A traveled the distance in 2.4 h, and Car B traveled the distance in 4 h. Car A traveled 22 mph faster than Car B. How fast did Car A travel?

OpenStudy (cutestgirlaround):

@Awolflover1

OpenStudy (dumbcow):

distance = rate*time Let distance be D, speed of car A be "a" , car B be "b" Car A: ---> D = 2.4a Car B: --- > D = 4b By substitution you can say: 2.4a = 4b We are given that a is 22 greater than b ----> a = b+22 or b = a - 22 By substitution: 2.4a = 4(a - 22) Solve for a

OpenStudy (cutestgirlaround):

4

OpenStudy (cutestgirlaround):

@F_Jayyy

OpenStudy (cutestgirlaround):

@RandomeQuestions @Frozen15 @dumbcow

OpenStudy (cutestgirlaround):

@Lolipop_Drum312

OpenStudy (dumbcow):

what is your question? do you follow my explanation

OpenStudy (lolipop_drum312):

hm... sorry dear idk i'm not really a math wizard but i can look it up for you if ya want

OpenStudy (lolipop_drum312):

but i can tell ya this: first you need to break it down they gave us the time for each car

OpenStudy (lolipop_drum312):

car A = 2.4 h and car B = 4 h

OpenStudy (lolipop_drum312):

so you need to find the speed so let's call it x

OpenStudy (lolipop_drum312):

so the travel would of the car be that car A travelled 22mph faster than B

OpenStudy (lolipop_drum312):

of the car*

OpenStudy (lolipop_drum312):

which makes A= x+22 mph and car B = x

OpenStudy (lolipop_drum312):

@dumbcow you can corrrect me or find someone to correct me if i'm wrong

OpenStudy (lolipop_drum312):

your choice but for now i hoped this help

OpenStudy (lolipop_drum312):

@NarutoFan2016

OpenStudy (lolipop_drum312):

he's gonna help and see if i'm right dear :)

OpenStudy (cutestgirlaround):

ok thanks for your help

OpenStudy (lolipop_drum312):

did i get it right my answer was that the car A = x+22 mph and car B = x

OpenStudy (cutestgirlaround):

I dont know yet

OpenStudy (narutofan2016):

close to 30+22=52 mph

OpenStudy (lolipop_drum312):

so its actually 52 mph?

OpenStudy (cutestgirlaround):

thanks I thought around 50